OpenStudy (anonymous):

If f(x)=4x-1 and g(x)=3x-5, what is g(f(2)?

OpenStudy (anonymous):

\[f(2)=4\times 2-1=8-1=7\] is a start now you need \(g(7)\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow g(4(2) - 1)}\) \( \color{Black}{\Rightarrow g(8 - 1)}\) \( \color{Black}{\Rightarrow g(7)}\)

OpenStudy (anonymous):

f(2)=4(2)-1 =8-1 =7 g(f(2))=g(7)=3(7)-5 =21-5 =16